6. You receive a prescription written for a topical solution which calls for 45 g to be dispensed. The recipe for the solution is as follows: (How much of each ingredient is needed to produce 45 grams of topical solution?)
Betamethasone valerate 121 mg
Alcohol (95%) 20 ml
Propylene glycol 20 ml
Pluronic F127 22 g
Purified water qs 100 g
1
Expert's answer
2012-09-28T09:10:39-0400
Proportions for all ingridients is:
Betamethasone valerate 121/(121+20+20+22+100)=42.7%

Alcohol (95%) 20/(121+20+20+22+100)=7.07%

Propylene glycol 20/(121+20+20+22+100)=7.07%

Pluronic F127 22/(121+20+20+22+100)=7.78%

Purified water qs 100/(121+20+20+22+100)=35.3%so the amount needed for 45g is

Betamethasone valerate 42.7%*45=19.2

Alcohol (95%) 7.07%*45=3.2

Propylene glycol 7.07%*45=3.2

Pluronic F127 7.78%*45=3.5

Purified water qs 35.3%*45=15.9